Write down the differential equation for a damped harmonic oscillator
1
Expert's answer
2020-02-20T09:41:00-0500

The Newton's second law for damped oscillator has a form


"ma+cv+kx=0."

Thus, the equation of motion of a damped harmonic oscillator is as follows


"m\frac{d^2x(t)}{dt^2}+c\frac{dx(t)}{dt}+kx(t)=0,"

or


"\ddot{x}(t)+2\beta\dot{x}(t)+\omega_0^2x(t)=0."

Here


"2\beta=\frac{c}{m},\quad \omega_0^2=\frac{k}{m}."